# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

### v2

@@ IV-A1 Implementation details
In this study, we employ the popular dlib face detector [67]. The detected face regions are enlarged by a factor of 1.3 and subsequently resized to 224×224 pixels before being input into the model. The proposed framework is implemented on the PyTorch [68] platform, and the model is trained using Adam optimizer [69] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process. Our default architecture employs the ViT-B as the backbone, and the ViT-B parameters are initialized with ImageNet-21k [70] weights. We set the batch size to 48, as a single 2080Ti GPU can accommodate a maximum of 48 images at a resolution of $\times224$. During training, we use AUC as the validation metric. The training process consists of 30,000 iterations, as we observe that the validation AUC converges after about 20,000 iterations.
@@ IV-A3 Evaluation Metrics
IV-B Cross-manipulation evaluation
In real-world applications, face forgery detectors inevitably encounter data generated by new forgery techniques, leading to significant performance drops. To evaluate the generalization capability, we train the models on the training sets comprising three source manipulation types of FF++. Subsequently, we evaluate the trained models on the test sets of the unseen target manipulation type. Table I presents the detection AUC and ACC of different models on the FF++ high-quality (c23) dataset. We bold the best detection results and underline the second-best results. It can be observed that OSDFD achieves SOTA AUC and ACC performances across all experiments.
Moreover, generalization to poor-quality data is crucial for real-world face forgery detectors. Therefore, we conduct the cross-manipulation experiments on low-quality (c40) data, as shown in Table II, where OSDFD outperforms previous approaches by a significant margin. This underscores the proposed method’s outstanding generalizability across diverse data qualities. Additionally, our approach demonstrates substantial improvements over the ViT-B baseline, demonstrating the effectiveness of the designed framework. Furthermore, compared to our conference version SCLoRA [27], our method consistently achieves superior performance on both c23 and c40 data, which can be attributed to the effectiveness of the designed forgery style mixture module and local CDC feature extractor. Finally, OSDFD comprises the fewest 1.34M trainable parameters, facilitating its deployment in real-world applications, particularly on mobile devices.
TABLE III: Cross-dataset evaluation on six unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Methods highlighted in blue denote video-level results. The trainable parameter number of each model is reported to evaluate the training efficiency.
@@ IV-B Cross-manipulation evaluation
OSDFD Ours 1.34M 83.35 24.47 78.53 28.46 76.28 31.17 75.52 31.99 95.54 9.90 84.61 24.09 82.31 25.01
@@
OSDFD Ours 1.34M 90.81 16.94 81.42 24.90 79.49 28.31 78.74 29.37 98.89 4.48 90.01 17.50 86.56 20.25
@@ IV-C Cross-dataset evaluation
Due to the significant domain gaps between training and testing datasets [83] in face forgery detection, cross-dataset evaluation has been a persistent challenge. We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors. To ensure fair comparisons in cross-dataset experiments, we follow previous works [63, 76, 16] by using the FF++ (c23) dataset [22] as the training set. This dataset includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures. To balance positive and negative samples during training, real faces are augmented fourfold. Additionally, during the forgery style mixture stage, we randomly mix the four forgery types. We test the trained models on six popular Deepfake datasets, including CelebDF-v2 (CDF) [84], WildDeepfake (WDF) [25], DeepFake Detection Challenge (DFDC) [23], DeepFake Detection Challenge Preview (DFDC-P) [23], DeepForensics-1.0 (DFR) [24], and Face Forensics in the Wild (FFIW) [26]. These testing datasets simulate real-world scenarios well, encompassing various environmental variables. AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code. The upper table presents the frame-level detection results, while the blue-highlighted bottom table shows video-level results. Video-level detection scores are calculated by averaging all frame scores within each video. The best and second-best results are highlighted in bold and underlined, respectively. The trainable parameter number of each detector is provided in the table for reference. Our method, OSDFD, achieves the highest average detection performance in both frame-level and video-level evaluation, with only 1.34M activated parameters. Compared to the ViT-B backbone, OSDFD consistently demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability. Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style mixture module.
@@ IV-D Robustness evaluation
We incorporate five severity levels for each perturbation type to mimic unpredictable environments. The robustness evaluation results are depicted in Fig. 8. OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications. All models depicted in Fig. 8 are trained on the FF++ c23 training set, with distortions applied to the corresponding test set. As the severity of distortion increases, all face forgery detection methods experience performance degradation. Compared to previous arts, our proposed approach achieves superior or comparable AUC performance across all six types of perturbation. Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations. Our proposed style mixture module contributes to enhancing model robustness by increasing the diversity of forgery domains, thereby reducing the distribution gap between clean training data and corrupted testing data.
@@
TABLE IV: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
@@
IV-E Ablation study
@@
IV-E1 Effectiveness of the designed modules
We conduct ablation experiments to study the impact of each designed module. Table IV reports the average cross-manipulation (c23) detection AUC and ACC scores across four experimental trials. The designed LoRA and Adapter modules significantly enhance the model’s generalizability. The combination of LoRA and Adapter enables the model to simultaneously explore global and local anomalies within input forgery face images, enhancing the final detection performance. The forgery-style mixture module further boosts the AUC and ACC scores by augmenting the forgery source domains during training. The Single-Center Loss (SCL) encourages the real features to cluster more tightly while pushing features from fake images away from the real center. This design leads to a clearer decision boundary between real and fake faces, further enhancing the generalizability of the detection. Overall, the experimental results presented in Table IV underscore the effectiveness of the designed components.
The proposed method showcases its adaptability by being flexibly adapted to other transformer backbones. OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones. As shown in Table V, we apply our method OSDFD to various models, including ViT [17] and Swin [85]. The results demonstrate that OSDFD consistently enhances performance across all backbones, leading to significant improvements in AUC scores in the cross-manipulation setting. Notably, the highest performance gains are observed in ViT-L (+0.167 AVG AUC) and Swin-L (+0.112 AVG AUC), suggesting that OSDFD is particularly effective when integrated with larger backbones. While the average cross-manipulation performance improves with increasing model capacity, larger models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment. These findings indicate that OSDFD is a flexible and effective module for improving the generalizability of transformer-based face forgery detection models.
In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model. Specifically, we only optimize the Adapter and LoRA layers while keeping the other ViT parameters fixed with ImageNet weights. We compare the PEFT strategy with the full fine-tuning (FT) strategy to validate its effectiveness. Table VI presents the cross-manipulation (c23) AUC results, where DF, FF, FS, and NT denote the testing manipulation types. The models are trained on three manipulations and tested on the remaining one. The two experimental settings employ two training strategies while maintaining the same network architectures. The results in Table VI demonstrate that the PEFT approach outperforms the fully FT strategy across all experimental trials. Notably, the average AUC score of PEFT exhibits a substantial enhancement, increasing from 68.2% to 84.3%. Hence, it can be concluded that the improvements in face forgery detection performance benefit from the PEFT strategy adopted. Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
TABLE V: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
@@ IV-E1 Effectiveness of the designed modules
TABLE VI: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
@@
Figure 10: Cross-dataset evaluation results using different seeds: (a) AUC; (b) EER.
IV-E2 Effectiveness of FSM
To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table VII. The cross-dataset evaluation results with and without FSM are presented in Table VII, showing that the inclusion of FSM consistently improves face forgery detection performance. Notably, FSM significantly enhances the FNR, indicating that it effectively mitigates forgery domain gaps between training and testing sets. This, in turn, enhances the model’s ability to detect forgery faces from another point of view.
IV-E3 Effectiveness of CDC
To validate the effectiveness of the CDC operator, we conduct an ablation study by integrating various conventional convolutions with different kernel sizes. Table VIII presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively. The Second-order Difference Convolution (SDC) module, as proposed by Fei et al. [54], employs a second-order local anomaly learning mechanism to detect anomalies in local regions. These baseline results are obtained by replacing the CDC convolution in Fig. 4 with the corresponding operators. The results demonstrate that the CDC operator consistently outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
IV-E4 Impacts of different ratios of mixed features
Our proposed method utilizes a feature mixture module to augment forgery domains in the feature space. As illustrated in Fig. 9, the $x$-axis represents the ratio of the mixed forgery features among all forgery features. We analyze the impact of this ratio by plotting the cross-manipulation (c23) ACC scores against ratio values in Fig. 9. Similarly, we present the average score across four cross-manipulation evaluation trials. A ratio of ‘0.0’ indicates the pure use of mixed forgery features during training, while ‘1.0’ means no forgery-style mixture is applied. Notably, relying solely on original or mixed forgery features results in limited face forgery detection accuracy. In contrast, combining original and mixed forgery features enhances the model’s generalization capability by augmenting the source forgery domains. The model achieves the highest average ACC score when the ratio of mixed forgery features is 50%. This observation underscores the effectiveness of the designed forgery feature mixture module from another perspective.
IV-F Visualization results
IV-F1 T-SNE feature map visualization
TABLE VII: Impacts of Forgery Style Mixture (FSM).
@@ IV-F1 T-SNE feature map visualization
IV-F2 Grad-CAM map visualization

### v3

@@ IV-A1 Implementation details
In this study, we employ the popular dlib face detector [84]. The detected face regions are enlarged by a factor of 1.3 and subsequently resized to 224×224 pixels before being input into the model. The proposed framework is implemented on the PyTorch [85] platform, and the model is trained using Adam optimizer [86] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process. In the proposed OSDFD framework, the ViT-Base model pre-trained on ImageNet-21K [87] serves as the backbone. It consists of 12 transformer blocks, each comprising a 12-head self-attention layer and a multi-layer perceptron (MLP) layer. We use the ViT-L/14 as the CLIP backbone. We set the batch size to 48, and the loss weight $\lambda$ in Eq. (10) is set to 1. During training, we use AUC as the validation metric. The training process consists of 30,000 iterations, as we observe that the validation AUC converges after about 20,000 iterations.
@@ IV-A3 Evaluation Metrics
TABLE III: Cross-dataset evaluation on six unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Top and bottom tables show frame- and video-level results, respectively. The trainable parameter number of each model is also reported. We separate the CLIP-based methods from others.
@@
OSDFD (ViT-B) Ours 1.34M 83.35 24.47 78.53 28.46 76.28 31.17 75.52 31.99 95.54 9.90 84.61 24.09 82.31 25.01
ForAda* [49] CVPR25 7.54M 87.61 21.71 79.90 27.60 83.87 24.74 79.42 28.08 93.51 13.96 77.66 29.62 83.67 24.29
UDD* [97] AAAI25 0.66M 82.68 25.20 75.92 31.31 79.56 27.92 73.67 32.42 90.93 16.84 78.79 28.10 80.26 26.97
CLIP [83] ICML21 303.18M 81.23 26.46 76.41 31.55 77.97 29.76 72.54 33.97 95.16 11.69 83.63 25.19 81.15 24.77
OSDFD (CLIP) Ours 2.89M 88.32 20.45 85.39 21.97 79.42 30.00 77.88 30.15 99.17 3.46 93.71 14.88 87.32 20.15
@@
OSDFD (ViT-B) Ours 1.34M 90.81 16.94 81.42 24.90 79.49 28.31 78.74 29.37 98.89 4.48 90.01 17.50 86.56 20.25
ForAda* [49] CVPR25 7.54M 93.87 14.41 85.64 22.47 91.40 16.53 83.77 23.51 97.48 8.46 85.12 23.22 89.55 18.10
UDD* [97] AAAI25 0.66M 90.31 17.29 78.90 32.20 84.97 23.76 77.14 29.85 96.17 11.44 85.54 23.51 85.51 23.01
CLIP [83] ICML21 303.18M 88.87 21.47 80.33 28.54 80.17 27.76 76.07 30.94 98.67 6.97 90.91 17.22 85.83 22.15
OSDFD (CLIP) Ours 2.89M 93.23 16.29 87.16 19.51 83.13 25.76 80.65 27.53 99.62 1.49 97.20 9.76 90.17 16.72
IV-B Cross-manipulation evaluation
In real-world applications, face forgery detectors inevitably encounter data generated by new forgery techniques, leading to significant performance drops. To evaluate the generalization capability, we train the models on the training sets comprising three source manipulation types of FF++. Subsequently, we evaluate the trained models on the test sets of the unseen target manipulation type. Table I presents the detection AUC and ACC of different models on the FF++ high-quality (c23) dataset. We bold the best detection results and underline the second-best results. We apply OSDFD to two backbone networks, ViT-B and CLIP (ViT-L/14). The results show that OSDFD built on ViT-B surpasses previous methods, while OSDFD (CLIP) achieves SOTA AUC and ACC across all evaluation settings.
Moreover, generalization to poor-quality data is crucial for real-world face forgery detectors. Therefore, we conduct the cross-manipulation experiments on low-quality (c40) data, as shown in Table II, where OSDFD outperforms previous approaches by a clear margin under three settings. This underscores the proposed method’s outstanding generalizability across diverse data qualities. OSDFD exhibits relatively lower performance on the (DF, FS, NT $\rightarrow$ FF) setting under low-quality cross-manipulation evaluation. This may be because the PEFT framework and style mixture module are well suited for capturing appearance-related anomalies but are less effective at generalizing to manipulations dominated by expression and motion artifacts, thereby limiting transferability to FF. Additionally, our approach demonstrates substantial improvements over the ViT-B baseline, demonstrating the effectiveness of the designed framework. Furthermore, compared to our conference version SCLoRA [28], our method achieves superior performance on both c23 and c40 data, which can be attributed to the effectiveness of the designed forgery style mixture module and local CDC feature extractor. Moreover, OSDFD (CLIP) exhibits good generalizability on low-quality data while requiring only 2.89M trainable parameters, making it well suited for deployment in open-set scenarios.
TABLE IV: Cross-dataset evaluations on DF40 (CDF) and FakeAVCeleb datasets.
Method FS (CDF) FR (CDF) EFS (CDF) FakeAVCeleb Average
AUC EER AUC EER AUC EER AUC EER AUC EER
SBI[15] 70.01 35.20 48.90 50.98 65.61 39.21 85.87 23.07 67.59 37.11
CFM[79] 78.65 28.73 69.11 35.96 71.49 34.33 82.84 26.27 75.52 31.32
DCL[16] 72.67 35.04 62.88 41.01 63.18 41.20 87.64 21.00 71.59 34.56
LAA-Net[36] 70.46 34.99 55.24 46.91 67.07 38.07 88.20 19.74 70.24 34.92
DiffusionFake[82] 70.34 35.75 55.37 46.85 63.70 41.57 85.91 22.54 68.83 36.67
ViT-B [17] 67.78 37.19 55.65 46.57 60.50 42.64 82.98 26.14 66.72 38.13
OSDFD (ViT-B) 82.50 24.59 64.78 39.22 77.68 26.65 84.04 23.03 77.25 28.37
CLIP [83] 78.93 29.21 62.33 42.49 71.23 34.10 82.86 25.27 73.83 32.76
ForAda[49] 83.33 25.13 63.76 40.73 82.30 25.34 83.67 25.41 78.26 29.15
UDD[97] 61.44 41.64 59.01 43.63 52.64 47.54 90.72 16.02 65.95 37.20
OSDFD (CLIP) 90.77 16.97 71.06 35.00 90.09 18.70 92.95 14.54 86.21 21.30
TABLE V: Efficiency evaluations w/ and w/o OSDFD.
[…]
(98.44% $\downarrow$) (9.34% $\uparrow$) (22.49% $\downarrow$) (+1.31) (+0.16) (+8.37)
[…]
(99.05% $\downarrow$) (20.48% $\uparrow$) (7.91% $\downarrow$) (+3.68) (+0.71) (+6.17)
@@ IV-C Cross-dataset evaluation
Due to the significant domain gaps between training and testing datasets [100, 101] in face forgery detection, cross-dataset evaluation has been a persistent challenge. We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors. In cross-dataset evaluation, our method OSDFD adopts the FF++ (c23) dataset [23] as the training set for fair comparisons. Our training set includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures, and the four manipulation types are randomly mixed during the forgery style mixture. To balance positive and negative samples during training, real faces are augmented fourfold. We test the trained models on six popular Deepfake datasets, including CelebDF-v2 (CDF) [102], WildDeepfake (WDF) [26], DeepFake Detection Challenge (DFDC) [24], DeepFake Detection Challenge Preview (DFDC-P) [24], DeepForensics-1.0 (DFR) [25], and Face Forensics in the Wild (FFIW) [27]. These testing datasets simulate real-world scenarios well, encompassing various environmental variables. AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code. The upper table presents the frame-level detection results, while the bottom table shows video-level results. Video-level detection scores are calculated by averaging all frame scores within each video. Since CLIP serves as a more powerful backbone for face forgery detection, we separate the CLIP-based methods (ForAda [49], UDD [97], CLIP [83], and OSDFD (CLIP)) from the other approaches to ensure fair comparisons. The best and second-best results are highlighted in bold and underlined, respectively. The trainable parameter number of each detector is provided in the table for reference. OSDFD (ViT-B) achieves promising average detection performance in both frame-level and video-level evaluations. In contrast, OSDFD built on the CLIP backbone attains better AUC and EER, as CLIP’s multimodal pretraining enables it to learn representations that align visual and semantic concepts, thereby enhancing generalization to distribution shifts. OSDFD’s relatively lower performance on DFDC and DFDC-P can be attributed to their large domain gaps with FF++ and the suppression of forgery artifacts by compression, which pose significant challenges for PEFT-based methods. Compared to the ViT-B and CLIP backbones, OSDFD demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability. Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style-mixture module.
We expand our evaluation to two other recently released and challenging datasets: DF40 [103] and FakeAVCeleb [104]. DF40 encompasses 40 diverse face forgery methods. Its domain-specific fake data generated from FF++ and CDF real faces. FakeAVCeleb is a multimodal benchmark comprising 20,000 videos, built from real YouTube celebrity footage across four ethnic backgrounds. We train our models on FF++ and report frame-level cross-dataset evaluations in Table IV. To prevent facial identity leakage, we test on three challenging DF40 subsets (FS (CDF), FR (CDF), and EFS (CDF)) derived from CDF real faces, and additionally evaluate on FakeAVCeleb (visual modality only). The upper part of Table IV shows results with conventional backbones, while the lower part reports CLIP-based models. Despite the sophistication of the testing datasets, OSDFD (ViT-B) achieves the best average AUC in the upper table, while OSDFD (CLIP) consistently outperforms all compared methods across all subsets, demonstrating superior generalizability on recent challenging datasets. The superior performance of CLIP-based models compared to ViT-B arises from CLIP’s large-scale multimodal pretraining, which equips it with stronger generalization ability to unseen forgery patterns.
@@ IV-D Robustness evaluation
We incorporate five severity levels for each perturbation type to mimic unpredictable environments. The robustness evaluation results are depicted in Fig. 8. OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications. All models depicted in Fig. 8 are trained on the FF++ c23 training set, with distortions applied to the corresponding test set. As the severity of distortion increases, all face forgery detection methods experience performance degradation. The average robustness results across the six distortion types indicate that both OSDFD (ViT-B) and OSDFD (CLIP) exhibit superior resilience to common image perturbations. OSDFD (ViT) demonstrates a lower robustness than EfficientNet-B4 under brightness perturbations. A potential reason lies in our use of the PEFT strategy. Pretrained knowledge captures real visual statistics (e.g., natural lighting), whereas forgeries frequently violate these subtle priors (e.g., illumination inconsistencies), leaving detectable forgery cues. However, brightness perturbations during robustness evaluation can obscure these artifacts, thereby diminishing our method’s performance. The designed FSM module enhances robustness by generating broader feature distributions to mitigate overfitting, introducing variability that improves resistance to perturbations, and guiding the model to focus on intrinsic forgery cues rather than superficial styles.
@@
TABLE VI: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
@@
IV-E Efficiency evaluations
OSDFD demonstrates high training efficiency and strong generalization capability with only negligible additional FLOPs. By introducing a small number of trainable parameters, OSDFD achieves lower storage requirements, faster training, and reduced GPU memory consumption compared with the vanilla backbones.
We conducted comprehensive efficiency evaluations of OSDFD using two backbones, ViT-B and CLIP, as presented in Table V. All experiments were implemented on an NVIDIA A40 GPU and an Intel(R) Xeon(R) Gold 6530 CPU. The training speed and inference time were measured over five trials, while the training speed and GPU memory were recorded with a batch size of 32. For reference, we also report the average AUC (AVG. AUC), representing the mean performance across six unseen datasets.
Compared with the vanilla ViT-B and CLIP backbones, OSDFD demonstrates strong parameter efficiency by introducing only a small number of trainable parameters. This design results in lower storage requirements, as the backbone weights can be stored in the cloud while only minimal parameters are retained locally. Moreover, OSDFD enables faster training because fewer parameters are updated during optimization. As shown in Table V, OSDFD improves training speed by 9.34% and 20.48% over the ViT-B and CLIP baselines, respectively. This efficiency makes OSDFD well-suited for frequent model updates, such as continual learning with newly emerging forgery data in open-set Deepfake detection. Furthermore, OSDFD effectively reduces GPU memory consumption during training, making it practical for deployment on resource-constrained devices.
The designed PEFT and forgery-style-mixture components introduce only a marginal increase in model size, leading to negligible additional inference time and computational overhead. Specifically, OSDFD incurs only 1.31 ms and 3.68 ms inference time, and introduces merely 0.16 G and 0.71 G additional FLOPs for the ViT-B and CLIP backbones, respectively, which are acceptable for real-world applications.
Beyond training efficiency, OSDFD provides additional benefits. The forgery-aware PEFT and forgery-style-mixture modules effectively integrate forgery-specific knowledge while preserving the pretrained model’s general representations. This design mitigates overfitting to forgery data and enhances generalization performance.
IV-F Ablation study
@@
IV-F1 Effectiveness of the designed modules
We conduct ablation experiments to study the impact of each designed module. Table VI reports the average cross-manipulation (c23) detection AUC and ACC scores across four experimental trials. The designed LoRA and Adapter modules significantly enhance the model’s generalizability. The combination of LoRA and Adapter enables the model to simultaneously explore global and local anomalies within input forgery face images, enhancing the final detection performance. The forgery-style-mixture module further boosts the AUC and ACC scores by augmenting the forgery source domains during training. The Single-Center Loss (SCL) encourages the real features to cluster more tightly while pushing features from fake images away from the real center. This design leads to a clearer decision boundary between real and fake faces, further enhancing the generalizability of the detection. Overall, the experimental results presented in Table VI underscore the effectiveness of the designed components.
The proposed method showcases its adaptability by being flexibly adapted to other transformer-based backbones. OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style-mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones. As shown in Table VII, we apply our method OSDFD to various models, including ViT [17] and Swin [105]. The results demonstrate that OSDFD enhances performance across these backbones, leading to significant improvements in AUC scores in the cross-manipulation setting. Notably, the highest performance gains are observed in ViT-L (+0.167 AVG AUC) and Swin-L (+0.112 AVG AUC), suggesting that OSDFD is particularly effective when integrated with larger backbones. While the average cross-manipulation performance improves with increasing model capacity, larger models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment. These findings indicate that OSDFD is a flexible and effective module for improving the generalizability of transformer-based face forgery detection models.
TABLE VII: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
@@ IV-F1 Effectiveness of the designed modules
In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model. Specifically, we only optimize the Adapter and LoRA layers while keeping the other ViT parameters fixed with ImageNet weights. We compare the PEFT strategy with the full fine-tuning (FT) strategy to validate its effectiveness. Table VIII presents the cross-manipulation (c23) AUC results, where DF, FF, FS, and NT denote the testing manipulation types. The models are trained on three manipulations and tested on the remaining one. The two experimental settings employ two training strategies while maintaining the same network architectures. The results in Table VIII demonstrate that the PEFT approach outperforms the fully FT strategy across all experimental trials. Notably, the average AUC score of PEFT exhibits a substantial enhancement, increasing from 68.2% to 84.3%. Hence, it can be concluded that the improvements in face forgery detection performance benefit from the PEFT strategy adopted. Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
Although ImageNet is not face-specific, its learned representations of natural visual statistics (e.g., smooth textures, natural lighting) remain highly relevant, as forgeries often violate these subtle priors (e.g., abnormal texture smoothness, illumination mismatches). A backbone pretrained on ImageNet therefore provides strong priors regarding what “real” images should look like, facilitating more effective forgery detection. Retaining these pretrained weights preserves natural image priors that help distinguish real from fake distributions, while PEFT introduces forgery-aware adaptability without overwriting the pretrained knowledge, thereby improving generalization and robustness to unseen domains. Moreover, since forgery artifacts are diverse and continuously evolving techniques, a backbone grounded in natural image distributions offers a stable reference frame, enabling consistent detection across varied manipulation types.
TABLE VIII: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
@@
TABLE IX: Impacts of Forgery Style Mixture (FSM).
@@
IV-F2 Effectiveness of FSM
To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table IX. The cross-dataset evaluation results with and without FSM are presented in Table IX, showing that the inclusion of FSM effectively improves face forgery detection performance. Notably, FSM significantly enhances the FNR, indicating that it effectively mitigates forgery domain gaps between training and testing sets. This, in turn, enhances the model’s ability to detect forgery faces from another point of view.
IV-F3 Impacts of training forgery type number
To investigate the effect of the number of training forgery types, we conducted ablation experiments under the cross-dataset setting. As shown in Table X, without the Forgery Style Mixture (FSM) module, incorporating more forgery types improves cross-dataset performance, which aligns with the expectation that larger and more diverse training sets enhance generalizability. In contrast, when the FSM module is applied, the model achieves further performance gains on unseen datasets, further confirming its effectiveness in diversifying training forgery style statistics. Moreover, the effectiveness of FSM increases with the number of forgery types. These findings suggest a promising direction for incorporating additional forgery domains to further enrich forgery style diversity and enhance generalizability.
IV-F4 Effectiveness of CDC
To validate the effectiveness of the CDC operator, we conduct an ablation study by integrating various conventional convolutions with different kernel sizes. Table XI presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively. The Second-order Difference Convolution (SDC) module, as proposed by Fei et al. [70], employs a second-order local anomaly learning mechanism to detect anomalies in local regions. These baseline results are obtained by replacing the CDC convolution in Fig. 4 with the corresponding operators. The results demonstrate that the CDC operator outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
IV-F5 Impacts of different $\lambda$ values
To investigate the effect of the loss weight of $L_{SCL}$, we conducted ablation experiments under the cross-dataset setting, as reported in Table XII. Compared with the baseline ($\lambda$=0), incorporating the Single-Center Loss $L_{SCL}$ effectively enhances the model’s generalizability. We varied the $\lambda$ value from 0.6 to 1.4 and found that $\lambda$=1.0 achieves better performance across unseen datasets.
Figure 11: T-SNE visualizations on FF++ and CDF datasets. (a) ViT-B baseline; (b) ViT-B with Forgery Style Mixture (FSM) module. (DF: Deepfakes, $F$: Face2Face, FS: FaceSwap, NT: NeuralTextures, OR: Original.)
IV-F6 Impacts of different ratios of mixed features
Our proposed method utilizes a feature mixture module to augment forgery domains in the feature space. As illustrated in Fig. 9, the $x$-axis represents the ratio of the mixed forgery features among all forgery features. We analyze the impact of this ratio by plotting the cross-manipulation (c23) ACC scores against ratio values in Fig. 9. Similarly, we present the average score across four cross-manipulation evaluation trials. A ratio of ‘0.0’ indicates the pure use of mixed forgery features during training, while ‘1.0’ means no forgery-style-mixture is applied. Notably, relying solely on original or mixed forgery features results in limited face forgery detection accuracy. In contrast, combining original and mixed forgery features enhances the model’s generalization capability by augmenting the source forgery domains. The model achieves the highest average ACC score when the ratio of mixed forgery features is 50%. This observation underscores the effectiveness of the designed forgery feature mixture module from another perspective.
TABLE X: Ablation experiment on training manipulation type number (cross-dataset evaluation).
FSM DF $F$ FS NT CDF DFR WDF
AUC EER AUC EER AUC EER
- - - - 75.85 32.08 91.22 15.86 72.13 34.43
- - - 77.51 29.80 92.13 15.61 73.68 33.81
- - 78.03 29.71 92.44 14.78 74.26 31.91
- - 78.12 30.01 93.01 14.57 75.18 31.82
- 79.88 28.26 93.77 12.88 76.75 30.89
- 80.76 26.58 94.18 11.69 76.68 30.76
83.35 24.47 95.54 9.90 78.53 28.46
IV-G Visualization results
IV-G1 T-SNE feature map visualization
We visualize the embedding distributions of ViT-B baseline (a) w/o and (b) w/ the Forgery Style Mixture (FSM) module in Fig. 11 to demonstrate the effectiveness of FSM, which is a key component of our method. The models are trained on the FF++ training sets and tested on both the FF++ test sets and the unseen CDF dataset. Feature centers of each domain are marked with $\star$”. In Fig. 11(a), the FF++ (OR) and CDF (Real) domains are closely aligned, whereas the unseen CDF (Fake) domain is significantly distant from the FF++ fake domains. This explains why the model effectively detects real faces but struggles with fake ones. Motivated by this observation, we design a FSM module to diversify the fake training data and expand the fake feature space without introducing additional data. As shown in Fig. 11(b), the FSM module effectively enlarges the forgery feature space, leading to greater overlap between the FF++ fake domains and the unseen CDF (Fake) domain. Consequently, the decision boundary becomes more robust, enabling improved detection of unseen face forgeries.
IV-G2 Grad-CAM map visualization